OpenStudy (anonymous):

What is the solution to the system of equations? 2x = 10 - y -2y = 22 - 3x hi, im not the best in math and i need help with this equation.Can you explain to me how we would solve this problem?.

OpenStudy (zpupster):

2x = 10 - y -2y = 22 - 3x it is best to write the equations in this format Ax+By=C can you do that?

OpenStudy (anonymous):

yes I can

OpenStudy (zpupster):

do that and we can solve from there

OpenStudy (anonymous):

so I st it up and got the equation 2X-y=10 and 3X-2y+=22

OpenStudy (zpupster):

2x = 10 - y add y to both sides 2x+y=10 -2y = 22 - 3x and we are going to add 3x both sides 3x-2y=22

OpenStudy (anonymous):

why do we have to add 'y' to both sides?

OpenStudy (zpupster):

to get to the other side remember we are looking for this format Ax+By=C

OpenStudy (anonymous):

oh okay right

OpenStudy (anonymous):

did you delete your post from earlier? I cant seem to find it.

OpenStudy (anonymous):

its gone again?

OpenStudy (anonymous):

did you mean the original problem?

OpenStudy (anonymous):

2x = 10 - y -2y = 22 - 3x I copied and pasted it, this it what i got.

OpenStudy (zpupster):

2x+y=10 y=10-2x -2(10-2x) = 22 - 3x -20 +4x = 22-3x 7x = 42 x=6 2(6)+y=10 12+y=10 y=-2 check these i am sorry i made a mistake earlier

OpenStudy (anonymous):

okay i will

OpenStudy (texaschic101):

2x = 10 - y ---> 2x + y = 10 -2y = 22 - 3x ---> 3x - 2y = 22 now we have : 2x + y = 10 --->(2) 3x - 2y = 22 ------------ 4x + 2y = 20 -- result of multiplying by 2 3x - 2y = 22 ------------add 7x = 42 x = 6 now sub 6 in for x 2x + y = 10 2(6) + y = 10 12 + y = 10 y = 10 - 12 y = -2 check.. 2x + y = 10 2(6) - 2 = 10 12 - 2 = 10 10 = 10 (correct) I agree with zpupster :)

OpenStudy (anonymous):

Thanks you guys,I understand know :)

OpenStudy (texaschic101):

good to hear...it is probably easier to use the substitution method like zpupster did. I used the elimination method. This is just another way to do it. You have to decide which way is easier for you to understand, and then do it that way.

OpenStudy (anonymous):

yea true,thanks again,you were lots of help.

OpenStudy (texaschic101):

thats what we are here for...to help :)